# Minimax rates in variance and covariance changepoint testing

### v1

@@ 2 Results for univariate data
$\quad\ \ \ \left.\min(t_{0},n-t_{0})\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)^{2}=\rho\right\},$ (13)
@@
i.e. the sub-classes of $\mathcal{P}(w,u)$ in accordance with the null hypothesis and the alternative hypothesis, respectively. Let $\Psi$ denote the class of measurable functions $\psi:\mathbb{R}^{n}\mapsto\{0,1\}$. We define the minimax testing error $\mathcal{M}(\rho)$ by
$\mathcal{M}(\rho)=\mathcal{M}(\rho,n,\sigma,w,u)=\underset{\psi\in\Psi}{\inf}\left\{\underset{P\in\mathcal{P}_{0}(n,w,u,\sigma)}{\sup}\mathbb{E}_{P}\psi(X)+\underset{P\in\mathcal{P}(n,w,u,\rho)}{\sup}\mathbb{E}_{P}\left(1-\psi(X)\right)\right\},$ (17)
@@
Theorem 1 shows an interesting similarity between the change-in-variance problem and the change-in-mean problem for univariate data. The signal strength of a single change in mean from $\mu_{1}$ to $\mu_{2}$ at observation $t_{0}$ with noise level $\sigma^{2}$ is given by $\min(t_{0},n-t_{0})(\mu_{1}-\mu_{2})^{2}\sigma^{-2}$ (see Liu et al., 2021), and the minimax rate of testing is given by $\log\log(8n)$ for this problem as well.
@@
A noteworthy distinction between the change-in-variance and change-in-mean problems is the forms of the signal strengths. For the change-in-variance problem, the signal strength in (13) depends quadratically on a variance ratio for weak signals, and linearly for strong signals. Meanwhile, the signal strength for the change-in-mean problem does not exhibit such a phase transition. This discrepancy is intuitively explained by the sub-exponential tails of $X_{i}^{2}$ being heavier than the tails of $X_{i}$, the former only exhibiting sub-Gaussian behaviour near the center of the distribution.
@@ 2.1 Upper bound
However, the true changepoint location $t_{0}$ under the alternative is unknown, which motivates the application of the aforementioned testing procedure across a range of values of $t$. To attain the level of precision needed to reach the minimax testing rate, we use the geometrically increasing grid in Liu et al. (2021), given by
@@ 2.2 Lower bound
Proposition 2 is shown by bounding the Chi square divergence between two probability measures $\mathbb{P}_{0}$ and $\mathbb{P}$ on $\mathbb{R}^{n}$, where $\mathbb{P}_{0}$ is consistent with the null hypothesis and $\mathbb{P}$ is consistent with the alternative. Specifically, $\mathbb{P}_{0}$ is the distribution of $X=(X_{1},\ldots,X_{n})^{\top}$ when $X_{i}~{}\sim~{}\text{N}(0,\sigma^{2})$ independently for $i=1,\ldots,n$, and $\mathbb{P}$ is the distribution of $X$ induced by sampling the changepoint location $t_{0}$ uniformly from $\{2^{0},2^{1},\ldots,2^{\lfloor\log_{2}(n/2)\rfloor}\}$, and conditionally on $t_{0}$, sampling $X_{i}\sim\text{N}(0,\sigma^{2}-\kappa)$ independently for $i\leq t_{0}$ and $X_{i}\sim\text{N}(0,\sigma^{2})$ independently for $i>t_{0}$. The scaling factor $\kappa<\sigma^{2}$ is chosen so that $\text{Cov}(X|t_{0})\in\Theta(n,\rho)$ for all $t_{0}$.
@@ 3 Results for multivariate data
Now let $M(p,s)$ denote the space of $s$-sparse matrices in $\mathbb{R}^{p\times p}$. For any $t_{0}\in[n-1]$, $s\in[p]$ and $\rho>0$, we define $\Theta^{(t_{0})}(p,n,s,\sigma,\rho)$ to be the set given by
@@
$\left.\min(t_{0},n-t_{0})\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}=\rho\right\}.$ (29)
@@
$\mathcal{M}(p,n,s,u,\sigma,\rho)=\underset{\psi\in\Psi}{\inf}\left\{\underset{P\in\mathcal{P}_{0}(p,n,u,\sigma)}{\sup}\mathbb{E}_{P}\psi(X)+\underset{P\in\mathcal{P}(p,n,s,u,\sigma,\rho)}{\sup}\mathbb{E}_{P}\left(1-\psi(X)\right)\right\}.$ (33)
@@ 3.1 A lower bound on the minimax testing error
$\underset{v\in\mathcal{S}^{p-1}_{s}}{\sup}\ \frac{v^{\top}\Sigma_{1}v}{v^{\top}\Sigma_{2}v}\vee\frac{v^{\top}\Sigma_{2}v}{v^{\top}\Sigma_{1}v}$ $\geq 1+c\frac{\gamma}{\Delta}\vee\sqrt{c\frac{\gamma}{\Delta}},$ (34)
@@
for some absolute constant $c>0$ depending only on $\delta$, where $\gamma=\gamma(p,n,s)=s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)$ and $\Delta=\Delta(n,t_{0})=\min(t_{0},n-t_{0})$ is the effective sample size. The left hand side of (34) is the largest relative change in the variance of the data along some subspace of $\mathbb{R}^{p}$ spanned by an $s$-sparse vector. In the asymptotic regime where $n,s$ or $p$ diverge, the right hand side of (34) diverges if (and only if) $\Delta/\gamma\rightarrow 0$. If $\Delta$ is fixed, the right hand side of (34) grows at the rate of $\gamma=s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)$. In this case, the relative change in variance along some axis of variation must change by a factor of order $\gamma$ for changepoint detection to be possible, amounting to a very stringent requirement, especially when $s$ or $p$ become large. To see this, consider for instance the case where $s=p$, in which the necessary relative change must be of order at least $p$ for any test to successfully discriminate between $H_{0}$ and $H_{1}$ over the whole parameter space.
@@
The proof of Proposition 3 builds upon the techniques used in the proof of Proposition 2. The main strategy is to bound the Chi square divergence between two probability measures $\mathbb{P}_{0}$ and $\mathbb{P}$ on $\mathbb{R}^{pn}$, where $\mathbb{P}_{0}$ is consistent with the null hypothesis and $\mathbb{P}$ is consistent with the alternative. Specifically, $\mathbb{P}_{0}$ is the distribution of $X=(X_{1}^{\top},\ldots,X_{n}^{\top})^{\top}$ when $X_{i}\sim\text{N}_{p}(0,\sigma^{2}I)$ independently for all $i$. Moreover, $\mathbb{P}$ is the mixture distribution of $(X_{1},\ldots,X_{n})$, which conditional on the changepoint location $t_{0}$ and the pre-change covariance matrix $\Sigma$ satisfies $X_{i}\sim\text{N}_{p}(0,\Sigma)$ independently for $i\leq t_{0}$ and $X_{i}\sim\text{N}_{p}(0,\sigma^{2}I)$ independently for $i\geq t_{0}+1$. To generate $t_{0}$ and $\Sigma$, we sample $t_{0}$ uniformly from $\{2^{0},2^{1},\ldots,2^{\lfloor\log_{2}(n/2)\rfloor}\}$, and independently sample $\Sigma=\sigma^{2}I-\kappa uu^{\top}$, where $\kappa<\sigma^{2}$ is deterministic chosen to satisfy the signal strength condition, and $u$ is suitably sampled from the $s$-sparse unit sphere $S^{p-1}_{s}$. Here, the stochastic choice of $t_{0}$ contributes with the $\log\log(8n)$ term in the lower bound, while the stochastic choice of $u$ contributes with the $s\log(ep/s)$ term.
@@
Let us compare Proposition 3 to the minimax result in Wang et al. (2021). Their result concerns changepoint location estimation, which is a slightly different than the testing problem considered in this paper. Wang et al. assume that $\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}\leq\sigma^{2}/4$, where $\sigma^{2}$ denotes the maximum sub-Gaussian norm of the $X_{i}$. The latter assumption implicitly restricts the effective sample size to satisfy $\min(t_{0},n-t_{0})\gtrsim p$. Within this setup, they show that consistent estimation of a changepoint location is impossible as long as
@@
Under the assumption that $\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}\leq\sigma^{2}/4$, the left hand side of (35) agrees with the signal strength defined in (29), up to absolute constants. Thus, Proposition 3 significantly refines the minimax result in Wang et al. (2021) by accounting for the dependence on both the sparsity and the sample size.
@@
where $\left\lVert\mu_{1}-\mu_{2}\right\rVert_{0}=s$ is fixed and $\mu_{1},\mu_{2}\in\mathbb{R}^{p}$ and $t_{0}\in[n-1]$ are unknown. Given any $\delta\in(0,1)$, Liu et al. (2021) show that no test can discriminate between $H_{0,\text{mean}}$ and $H_{1,\text{mean}}$ with minimax testing error less than $\delta$ whenever
@@
$\Delta\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}$ $\lesssim s\log\left(\frac{ep}{s}\right)\vee\log\log(8n),$ (39)
where $\Delta$ is defined as above. Interestingly, the minimax testing rate for a change in the mean features a phase transition between the dense and sparse regimes, resulting in an elbow effect in the minimax rate in (38). Such a phase transition is not present for the covariance changepoint problem, in which no phase transition occurs with respect to the sparsity $s$. A different phase transition occurs, however, in the effect of the variance ratio in the left hand side of (39), being linear for large values of the variance ratio and quadratic for smaller values.
@@ 3.2 Upper bound on the minimax testing error
The statistic $S_{t,s}$ measures the absolute difference between the empirical covariance matrices of $X_{1},\ldots,X_{t}$ and $X_{n-t+1},\ldots,X_{n}$ in terms of the largest absolute $s$-sparse eigenvalue. When $s=p$, the statistic $S_{t,s}$ is reminiscent of the Covariance CUSUM statistic defined in Wang et al. (2021). Due to the sub-Gaussianity of the $X_{i}$, high-probability control over the tails of $S_{t,s}$ can be obtained under the null hypothesis using standard concentration inequalities for empirical covariance matrices. Meanwhile, under the alternative hypothesis, $S_{t,s}$ grows linearly with the signal strength, as both $\widehat{\Sigma}_{1,t}$ and $\widehat{\Sigma}_{2,t}$ are unbiased estimators of the pre- and post-change covariances, respectively. A natural testing procedure is therefore to reject the null hypothesis for large values of $S_{t,s}$. To guarantee statistical power over all possible changepoint locations, we apply $S_{t,s}$ over the grid $\mathcal{T}$ given in (21).
@@
The following theorem gives the theoretical performance of the test in Equation (42).
@@ 3.3 Adaptivity to noise level and sparsity
The testing procedure in (42) is minimax rate optimal in a large region of the parameter space, but requires knowledge of the sparsity $s$ and variance parameter $\sigma^{2}$. In this section we present a modified test statistic that is adaptive to these quantities. Since the test in (42) is only guaranteed to detect an $s$-sparse change in covariance when the effective sample size is of the same order as $\gamma=\gamma(p,n,s)=s\log(ep/s)\vee\log\log(8n)$, a natural estimate of the (scaled) variance of $S_{t,s}$ is given by
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}^{(1)})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}^{(2)}),$ (47)
where $\widehat{\Sigma}^{(i)}=\widehat{\Sigma}_{i,\lceil\gamma(p,n,s)\rceil}$, the latter being defined in Equation (41). Here, the first and last $\lceil\gamma(p,n,s)\rceil$ data points contribute to the variance estimate, as no changepoints in these segments can be guaranteed to be detected by our testing procedure anyway. Given a tuning parameter $\lambda>0$, our adaptive test is defined by
@@
is a geometric grid of candidate sparsities. Note that this geometric grid is sufficient to retain power over all possible sparsities $s\in[p]$. Unlike the test in (42), the test in (48) is adaptive to the noise level by using an estimate for the noise level, and is adaptive to the sparsity level by testing over all sparsities in the grid $\mathcal{S}$.
@@
Note that the theoretical performance of the adaptive testing procedure in (48) is stronger than that of the testing procedure in (42), since the condition in Equation (51) is weaker than the condition in Equation (45). This is due to the adaptivity of (48) to both the sparsity and noise level of the data.

### v2

@@ 2 Results for univariate data
$\quad\ \ \ \left.\min(t_{0},n-t_{0})\left[\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)^{2}\right]=\rho\right\},$ (13)
@@
i.e. the sub-classes of $\mathcal{P}(w,u)$ in accordance with the null hypothesis and the alternative hypothesis, respectively. Let $\Psi$ denote the class of measurable functions $\psi:\mathbb{R}^{n}\mapsto\{0,1\}$. We define the minimax testing error $\mathcal{M}(\rho)=\mathcal{M}(\rho,n,\sigma,w,u)$ by
$\mathcal{M}(\rho)=\underset{\psi\in\Psi}{\inf}\left\{\underset{P\in\mathcal{P}_{0}(n,w,u,\sigma)}{\sup}\mathbb{E}_{P}\psi(X)+\underset{P\in\mathcal{P}(n,w,u,\rho)}{\sup}\mathbb{E}_{P}\left(1-\psi(X)\right)\right\},$ (17)
@@
Theorem 1 shows an interesting similarity between the change-in-variance problem and the change-in-mean problem for univariate data. The signal strength of a single change in mean from $\mu_{1}$ to $\mu_{2}$ at observation $t_{0}$ with noise level $\sigma^{2}$ is given by $\min(t_{0},n-t_{0})(\mu_{1}-\mu_{2})^{2}\sigma^{-2}$ (see Liu et al.,, 2021), and the minimax rate of testing is given by $\log\log(8n)$ for this problem as well.
@@
A noteworthy distinction between the change-in-variance and change-in-mean problems is the forms of the signal strengths. For the change-in-variance problem, the signal strength in (13) depends quadratically on a variance ratio for weak signals, and linearly for strong signals. Meanwhile, the signal strength for the change-in-mean problem does not exhibit such a phase transition. This discrepancy is intuitively explained by the sub-exponential tails of $X_{i}^{2}$ being heavier than the tails of $X_{i}$, the former only exhibiting sub-Gaussian behavior near the center of the distribution.
@@ 2.1 Upper bound
However, the true changepoint location $t_{0}$ under the alternative is unknown, which motivates the application of the aforementioned testing procedure across a range of values of $t$. To attain the level of precision needed to reach the minimax testing rate, we use the geometrically increasing grid in Liu et al., (2021), given by
@@ 2.2 Lower bound
The proof of Proposition 2 adopts techniques from Liu et al., (2021), and the proposition is shown by bounding the Chi square divergence between two probability measures $\mathbb{P}_{0}$ and $\mathbb{P}$ on $\mathbb{R}^{n}$, where $\mathbb{P}_{0}$ is consistent with the null hypothesis and $\mathbb{P}$ is consistent with the alternative. Specifically, $\mathbb{P}_{0}$ is the distribution of $X=(X_{1},\ldots,X_{n})^{\top}$ when $X_{i}~{}\sim~{}\text{N}(0,\sigma^{2})$ independently for $i=1,\ldots,n$, and $\mathbb{P}$ is the distribution of $X$ induced by sampling the changepoint location $t_{0}$ uniformly from $\{2^{0},2^{1},\ldots,2^{\lfloor\log_{2}(n/2)\rfloor}\}$, and conditionally on $t_{0}$, sampling $X_{i}\sim\text{N}(0,\sigma^{2}-\kappa)$ independently for $i\leq t_{0}$ and $X_{i}\sim\text{N}(0,\sigma^{2})$ independently for $i>t_{0}$. The scaling factor $\kappa<\sigma^{2}$ is chosen so that $\text{Cov}(X|t_{0})\in\Theta(n,\rho)$ for all $t_{0}$.
@@ 3 Results for multivariate data
Now, let $M(p,s)$ denote the space of $s$-sparse matrices in $\mathbb{R}^{p\times p}$. For any $t_{0}\in[n-1]$, $s\in[p]$ and $\rho>0$, we define $\Theta^{(t_{0})}(p,n,s,\sigma,\rho)$ to be the set given by
@@
$\left.\min(t_{0},n-t_{0})\left[\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}\right]=\rho\right\}.$ (29)
@@
$\mathcal{M}(\rho)=\underset{\psi\in\Psi}{\inf}\left\{\underset{P\in\mathcal{P}_{0}(p,n,u,\sigma)}{\sup}\mathbb{E}_{P}\psi(X)+\underset{P\in\mathcal{P}(p,n,s,u,\sigma,\rho)}{\sup}\mathbb{E}_{P}\left(1-\psi(X)\right)\right\}.$ (33)
@@ 3.1 A lower bound on the minimax testing error
$\underset{v\in S^{p-1}_{s}}{\sup}\ \frac{v^{\top}\Sigma_{1}v}{v^{\top}\Sigma_{2}v}\vee\frac{v^{\top}\Sigma_{2}v}{v^{\top}\Sigma_{1}v}$ $\geq 1+c\frac{\gamma}{\Delta}\vee\sqrt{c\frac{\gamma}{\Delta}},$ (34)
@@
for some absolute constant $c>0$ depending only on $\delta$, where $\gamma=\gamma(p,n,s)=s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)$ and $\Delta=\Delta(n,t_{0})=\min(t_{0},n-t_{0})$ is the effective sample size. The left-hand side of (34) is the largest relative change in the variance of the data along some subspace of $\mathbb{R}^{p}$ spanned by an $s$-sparse vector. In the asymptotic regime where $n,s$ or $p$ diverge, the right-hand side of (34) diverges if (and only if) $\Delta/\gamma\rightarrow 0$. If $\Delta$ is fixed, the right-hand side of (34) grows at the rate of $\gamma=s\log\left(\frac{ep}{s}\right)\vee\log\log(8n)$. In this case, the relative change in variance along some axis of variation must change by a factor of order $\gamma$ for changepoint detection to be possible, amounting to a very stringent requirement, especially when $s$ or $p$ become large. To see this, consider for instance the case where $s=p$, in which the necessary relative change must be of order at least $p$ for any test to successfully discriminate between $H_{0}$ and $H_{1}$ over the whole parameter space.
@@
The proof of Proposition 3 is similar to that of Proposition 2. The main strategy is to bound the Chi square divergence between two probability measures $\mathbb{P}_{0}$ and $\mathbb{P}$ on $\mathbb{R}^{pn}$, where $\mathbb{P}_{0}$ is consistent with the null hypothesis and $\mathbb{P}$ is consistent with the alternative. Specifically, $\mathbb{P}_{0}$ is the distribution of $X=(X_{1}^{\top},\ldots,X_{n}^{\top})^{\top}$ when $X_{i}\sim\text{N}_{p}(0,\sigma^{2}I)$ independently for all $i$. Moreover, $\mathbb{P}$ is the mixture distribution of $(X_{1},\ldots,X_{n})$, which conditional on the changepoint location $t_{0}$ and the pre-change covariance matrix $\Sigma$ satisfies $X_{i}\sim\text{N}_{p}(0,\Sigma)$ independently for $i\leq t_{0}$ and $X_{i}\sim\text{N}_{p}(0,\sigma^{2}I)$ independently for $i\geq t_{0}+1$. To generate $t_{0}$ and $\Sigma$, we sample $t_{0}$ uniformly from $\{2^{0},2^{1},\ldots,2^{\lfloor\log_{2}(n/2)\rfloor}\}$, and independently sample $\Sigma=\sigma^{2}I-\kappa uu^{\top}$, where $\kappa<\sigma^{2}$ is deterministic chosen to satisfy the signal strength condition, and $u$ is suitably sampled from the $s$-sparse unit sphere $S^{p-1}_{s}$. Here, the stochastic choice of $t_{0}$ contributes with the $\log\log(8n)$ term in the lower bound, while the stochastic choice of $u$ contributes with the $s\log(ep/s)$ term.
@@
Let us compare Proposition 3 to the minimax result in Wang et al., (2021). Their result concerns changepoint location estimation, which is a slightly different from the testing problem considered in this paper. Wang et al., assume that $\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}\leq\sigma^{2}/4$, where $\sigma^{2}$ denotes the maximum sub-Gaussian norm of the $X_{i}$. The latter assumption implicitly restricts the effective sample size to satisfy $\min(t_{0},n-t_{0})\gtrsim p$. Within this setup, they show that consistent estimation of a changepoint location is impossible as long as
@@
Under the assumption that $\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}\leq\sigma^{2}/4$, the left-hand side of (35) agrees with the signal strength defined in (29), up to absolute constants. Thus, Proposition 3 significantly refines the minimax result in Wang et al., (2021) by accounting for the dependence on both the sparsity and the sample size.
@@
where $\left\lVert\mu_{1}-\mu_{2}\right\rVert_{0}=s$ is fixed and $\mu_{1},\mu_{2}\in\mathbb{R}^{p}$ and $t_{0}\in[n-1]$ are unknown. Given any $\delta\in(0,1)$, Liu et al., (2021) show that no test can discriminate between $H_{0,\text{mean}}$ and $H_{1,\text{mean}}$ with minimax testing error less than $\delta$ whenever
@@
$\Delta\left\{\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}\right\}$ (39)
$\lesssim$ $s\log\left(\frac{ep}{s}\right)\vee\log\log(8n),$ (40)
where $\Delta$ is defined as above. Interestingly, the minimax testing rate for a change in the mean features a phase transition between the dense and sparse regimes, resulting in an elbow effect in the minimax rate in (38). Such a phase transition is not present for the covariance changepoint problem, in which no phase transition occurs with respect to the sparsity $s$. A different phase transition occurs, however, in the effect of the variance ratio in the left-hand side of (40), being linear for large values of the variance ratio and quadratic for smaller values.
@@ 3.2 Upper bound on the minimax testing error
The statistic $S_{t,s}$ measures the absolute difference between the empirical covariance matrices of $X_{1},\ldots,X_{t}$ and $X_{n-t+1},\ldots,X_{n}$ in terms of the largest absolute $s$-sparse eigenvalue. When $s=p$, the statistic $S_{t,s}$ is reminiscent of the Covariance CUSUM statistic defined in Wang et al., (2021). Due to the sub-Gaussianity of the $X_{i}$, high-probability control over the tails of $S_{t,s}$ can be obtained under the null hypothesis using standard concentration inequalities for empirical covariance matrices. Meanwhile, under the alternative hypothesis, $S_{t,s}$ grows linearly with the signal strength, as both $\widehat{\Sigma}_{1,t}$ and $\widehat{\Sigma}_{2,t}$ are unbiased estimators of the pre- and post-change covariances, respectively. A natural testing procedure is therefore to reject the null hypothesis for large values of $S_{t,s}$. To guarantee statistical power over all possible changepoint locations, we apply $S_{t,s}$ over the grid $\mathcal{T}$ given in (21).
@@
The following theorem gives the theoretical performance of the test in (43).
@@ 3.3 Adaptivity to noise level and sparsity
The testing procedure in (43) is minimax rate optimal in a large region of the parameter space, but requires knowledge of the sparsity $s$ and variance parameter $\sigma^{2}$. In this section we present a modified test statistic that is adaptive to these quantities. Since the test in (43) is only guaranteed to detect an $s$-sparse change in covariance when the effective sample size is of the same order as $\gamma(s)=\gamma(p,n,s)=s\log(ep/s)\vee\log\log(8n)$, a natural estimate of the (scaled) variance of $S_{t,s}$ is given by
$\widehat{\sigma}^{2}_{s}=\lambda_{\max}^{s}(\widehat{\Sigma}_{1,\lceil\gamma(s)\rceil})\wedge\lambda_{\max}^{s}(\widehat{\Sigma}_{2,\lceil\gamma(s)\rceil}),$ (48)
where $\widehat{\Sigma}_{i,t}$ is defined in (42), for $i=1,2$ and $t\in[n]$. In (48), the first and last $\lceil\gamma(p,n,s)\rceil$ data points contribute to the variance estimate, as no changepoints in these segments can be guaranteed to be detected by our testing procedure anyway. It is therefore in (48) implicitly assumed that $n\geq\gamma(s)$. Given a tuning parameter $\lambda>0$, our adaptive variant of the test in (43) is defined by
@@
is a geometric grid of candidate sparsities. Note that this geometric grid is sufficient to retain power for all possible sparsities $s\in[p]$. Unlike the test in (43), the test in (49) is adaptive to the noise level by using an estimate for the noise level, and is adaptive to the sparsity level by testing over all sparsities in the grid $\mathcal{S}$.
@@
Note that the theoretical performance of the adaptive testing procedure in (49) is stronger than that of the testing procedure in (43), since the condition in (52) is weaker than the condition in (46). This is due to the adaptivity of (49) to both the sparsity and noise level of the data.